The following is a report from a BLS survey taker: There were 65 people in the houses I visited. Ten of them were children under 16, 25 people had full-time jobs, and 5 had part-time jobs. There were 10 retirees, 5 full-time homemakers, 5 full-time students over age 16, and 2 people who were disabled and cannot work. The remaining people did not have jobs, but all said they would like one. One of these people had not looked actively for work for 3 months, however. Find the labor force, the unemployment rate, and the participation rate implied by the survey takers report. Labor force:32people Instructions: Enter your responses rounded to two decimal places. Unemployment rate: [ 625]0% Participation rate: -1%
Answer #1

Answer : Here children who are under 16 years = 10

Total population = 65 people

Working age population = 65 - 10 = 55 people

Participation rate = (Labor force / Working age population) * 100

=> Participation rate = (32 / 55) * 100 = 58.18%

Therefore, here the participation rate is 58.18% .